OpenStudy (anonymous):

what is the square for x^2-5x=10

OpenStudy (anonymous):

do u know the answer for this ?

OpenStudy (ddcamp):

The roots? First put it into standard form, \[x^2-5x=10 \rightarrow x^2-5x-10 = 0\] This isn't factor-able, so use the quadratic formula: \[\frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] With a = 1, b = -5, c =-10 \[\frac{ 5 \pm \sqrt{5^2 - 4(1)(-10)} }{ 2(1) }\rightarrow \frac{ 5\pm \sqrt{65} }{ 2 }\] Which doesn't simplify nicely, but is about 6.5311 and -1.5311